# Chapter 0 Hadron Production Processes

### v1

@@ 1 Single Pseudoscalar Meson Photoproduction
Single pseudoscalar meson photoproduction involves the interaction of a photon with a free proton, a bound neutron, or a whole nucleus. For studies of the baryon spectrum, one is normally interested in the first two of these. So a spin-1 particle (the photon, two helicity states) and a spin-$\frac{1}{2}$ particle (the nucleon) react to give a spin-0 particle (the pseudoscalar meson) and a spin-$\frac{1}{2}$ particle (the recoiling baryon). This gives eight spin combinations, of which four are possible within the parity-conserving strong interaction that has taken place. EM interaction does not conserve isospin, so multipole amplitudes contain isoscalar and isovector contributions of EM current.
@@
The four combinations are represented as amplitudes (Eqs. (21) and (22)), the exact form of which is a matter of choice. A theory of pion photoproduction was constructed in the 1950s. Kroll and Ruderman [125] were the first to derive model-independent predictions in the threshold region, a so-called low energy theorem (LET), by applying gauge and Lorentz invariance to the reaction $\gamma N\to\pi N$. The general formalism for this process was developed by Chew, Goldberger, Low, and Nambu (Geoffrey Foucar Chew (1924-2018), Marvin Leonard “Murph” Goldberger (1922-2014), Francis Eugene Low (1921-2007), and Yoichiro Nambu (1921-2015) [^22] Nambu was awarded half of the Nobel Prize in Physics in 2008 for the discovery in 1960 of the mechanism of spontaneous broken symmetry in subatomic physics, related at first to the strong interaction’s chiral symmetry and later to the electroweak interaction and Higgs mechanism.). The results are known as CGLN amplitudes [126] and a few years later helicity amplitudes [127] were derived. About a decade later, Berends, Donnachie, and Weaver analyzed the existing data in terms of a multipole decomposition and extracted the various multipole amplitudes contributing in a region up to an excitation energy of $500~{}\mathrm{MeV}$ [128]. These amplitudes are vital inputs to low-energy descriptions of hadron physics based on the chiral perturbation theory (ChPT) [129]. Within any of these bases, there are 16 possible bilinear combinations that are referred to as the “observables.”
@@
Proton data alone does not allow separation of isoscalar and isovector components [130].
@@
Accurate evaluation of EM couplings $N^{\ast}\to\gamma N$ and $\Delta\to\gamma N$ from meson photoproduction data remains a paramount task in hadron physics. Only with good data on both proton and neutron targets can one hope to disentangle isoscalar and isovector EM couplings of various $N^{\ast}$ and $\Delta^{\ast}$ resonances [131, 132], as well as isospin properties of non-resonant background amplitudes.
@@
The lack of $\gamma n\to\pi^{-}p$ and $\gamma n\to\pi^{0}n$ data [^23] Apart from lower-energy ($<700~{}\mathrm{MeV}$), there are data for the inverse $\pi^{-}$ photoproduction reaction, $\pi^{-}p\to\gamma n$ [99]. This process is free from complications associated with a deuteron target. However, there is a major disadvantage of using $\pi^{-}p\to\gamma n$: there is a large background from $\pi^{-}p\to\pi^{0}n\to\gamma\gamma n$, whose cross section is 5 to 500 times larger than $\pi^{-}p\to\gamma n$. does not allow us to be as confident about the determination of neutron couplings relative to those of protons.
@@
Since the neutron targets do not exist, it remains to use nuclear ones. In this case, when extracting information on the elementary reaction on the bound neutron from nuclear data, one should take into account the nuclear-medium effects, i.e., the final-state interaction (FSI) and Fermi-motion effects [133, 134] [^24] It is impossible to measure FSI experimentally [135]. Obviously, in the case of polarized measurements the FSI corrections are small or consistent with experimental uncertainties.
@@
Measurements of pion photoproduction on both proton and quasi-free neutron targets have a very long history, starting about 70 years ago with the discovery of the pion by the University of Bristol group [4]. Two years later, at the 1949 Spring Meeting of the US National Academy of Sciences, a preliminary account was given of some observations of mesons produced by the $335-\mathrm{MeV}$ photon beam from the Berkeley synchrotron [136] [^25] Edwin Mattison McMillan (1907-1991) shared the 1951 Nobel Prize in Chemistry with Glenn Theodore Seaborg (1912-1999), credited with being the first to produce a transuranium element, neptunium. In addition, McMillan co- invented the synchrotron with Vladimir Iosifovich Veksler (1907-1966). Starting with the use of bremsstrahlung facilities, pioneering results for $\gamma p\to\pi^{+}n$ (Fig. 14[136] and for $\gamma p\to\pi^{0}p$ [137] were obtained. One can possibly understand that the pion family is a triplet. Finding a neutral pion is much more difficult since it does not leave marks in photoemulsion or Wilson chambers (named after Charles Wilson).
@@
In addition to the free proton reactions as $\gamma p\to n\pi^{+}$, the experiment at Berkeley has utilized the loosely bound neutron $\gamma n\to p\pi^{-}$ at the photon energy $318\pm 10~{}\mathrm{MeV}$ [138]. So, the negative pion photoproduction experiment was done in the next 3 years after the positive one.
@@
Despite all the shortcomings of the first measurements (such as large normalization uncertainties, wide energy and angular binning, limited angular coverage, and so on), these data were crucial for the discovery of the first excited Delta and nucleon states using PWA for $\pi N$ elastic scattering data. The $\Delta(1232)3/2^{+}$ was determined by Enrico Fermi’s group [139]. While the second one $N(1440)1/2^{+}$, called “Roper,” (named after L. David Roper (born 1935) who discovered this baryonic state) came in several years after the $\Delta$-isobar [140] [^26] As an anecdote, it is worth mentioning that Lev Davidovich Landau, when becoming aware of Fermi’s discovery of the Delta resonance, did not believe in the existence of such a state, arguing A width of $120~{}\mathrm{MeV}$ what is that? The pion will make a quarter of the circle around the nucleon and that is supposed to be a pion-nucleon bound state? The concept of a meson-nucleon resonance was waiting to be established $\ldots$..
Figure 15: Database for $\gamma p\to\pi^{0}p$ (left) and $\gamma p\to\pi^{+}n$ (right). Experimental data from the SAID database [99] selected for 1996 through 2018. Right: Amount of data as a function of time. Full SAID database. The data shown as stacked histogram. Light shaded – cross sections, dark shaded – polarization data. The figure is adapted from Ref. [141].
Figure 16: Database for $\gamma n\to\pi^{-}p$ (left) and $\gamma n\to\pi^{0}n$ (right). Experimental data from the SAID database [99] selected for 1996 through 2018. Right: Amount of data as a function of time. Full SAID database. The data shown as stacked histogram. Light shaded – cross sections, dark shaded – polarization data. The figure is adapted from Ref. [141].
@@
The “neutron” database is significantly smaller than the “proton” one. The majority of single pseudoscalar meson photoproduction of the “neutron” target came from EM facilities at BNL and JLab, USA; MAMI, Germany; GRAAL, France (Fig. 16[141].
Studies of the $\gamma n\to\pi^{-}p$ and $\gamma n\to\pi^{0}n$ reactions can be carried out in quasi-free kinematics with deuteron targets. The reactions $\gamma d\to\pi^{-}p(p)$ and $\gamma d\to\pi^{0}n(p)$ in these kinematics have a fast, knocked-out nucleon and a slow proton spectator, and the slow proton is assumed not to be involved in the pion production process. In this quasi-free region, the reaction mechanism corresponds to the “dominant” impulse approximation (IA) diagram in Fig. 17(a) with the slow proton emerging from the deuteron vertex. Here, the differential cross section on the deuteron can be related to that on the neutron target in a well-understood way, see Ref. [142] and references therein. Fig. 17 illustrates this dominant IA diagram, as well as the leading terms of FSI corrections.
An energy and angle dependent FSI correction factor, R(E, $\theta$), can be defined as the ratio between the sum of three dominant diagrams in Fig. 17 and IA (the first of the diagrams). This can then be applied to the experimental $\gamma d$ data to get a two-body cross section for $\gamma n\to\pi^{-}p$ and $\gamma n\to\pi^{0}n$.
The GWU-ITEP FSI calculations (see Ref. [142] and references therein) are available over a broad energy range (threshold to $E=2.7~{}\mathrm{GeV}$), and for the full CM angular range ($\theta=0^{\circ}$ to $180^{\circ}$). Overall, the FSI correction factor $R<1.0$, while its value varies from 0.70 to 0.90 depending on the kinematics. The behavior of R is very smooth vs. pion production angle. There is a sizable FSI effect from the $S$-wave part of $pp$-FSI at small angles.
@@
R(E, $\theta$) is used as the FSI correction factor for the CLAS quasi-free $\gamma d\to\pi pN$ cross section averaged over the laboratory photon energy bin width. Note that the FSI correction grows rapidly to the forward direction ($\theta<30^{\circ}$). There are currently few measurements in this regime, so the uncertainty due to FSI for this reaction at forward angles does not cause too much concern. The contribution of uncertainty in FSI calculations to the overall systematic normalization uncertainty is estimated to be about 2%–3% (the sensitivity to the deuteron wave function is 1% and to the number of steps in the integration of the five-fold integrals is 2%). For the CLAS measurements, no sensitivity was found to the value of proton momentum used to determine whether or not it is a spectator.
@@
Eta and Eta’ Photoproduction on the Proton
@@
The experimental activity on the pseudoscalar $\eta^{(}548)$ and $\eta^{\prime}(958)$ mesons has a short history, not to the least because they are charge-neutral particles and, as such, are notoriously hard to detect. The singlet $eta^{\prime}$ meson, in addition, is much heavier than the octet part of the pseudoscalar nonet. Compared to the pion case, the data base is much smaller, as is evident from (Fig. 18).
@@
Figure 18: Database for $\gamma p\to\eta p$ (left) and $\gamma p\to\eta^{\prime}p$ (right). Experimental data from the SAID database [99] selected for 1996 through 2018. Right: Amount of data as a function of time. Full SAID database. The data shown as stacked histogram. Light shaded – cross sections, dark shaded – polarization data. The figure is adapted from Ref. [141].
Figure 19: Database for the Kaon photoproduction data on both proton and neutron targets. Experimental data from the SAID database [99] selected for 1996 through 2018. Right: Amount of data as a function of time. Full SAID database. The data shown as stacked histogram. Light shaded – cross sections, dark shaded – polarization data. The figure is adapted from Ref. [141].
Kaon Photoproduction

### v2

@@ 2 Single Pion Photoproduction
Single pseudoscalar meson photoproduction involves the interaction of a photon with a free proton, a bound neutron, or an entire nucleus. The growth of the database on photoproduction of $\pi^{0,+}$ meson production on the proton is depicted in Fig. 15. For studies of the baryon spectrum, one is normally interested in the first two of these. Thus, a spin-1 particle (the photon, two helicity states) and a spin-$\frac{1}{2}$ particle (the nucleon) react to produce a spin-0 particle (the pseudoscalar meson) and a spin-$\frac{1}{2}$ particle (the recoiling baryon). This yields eight spin combinations, of which four are possible within the parity-conserving strong interaction that has occurred. EM interaction does not conserve isospin, so multipole amplitudes contain isoscalar and isovector contributions of the EM current.
@@
The four combinations are represented as amplitudes (Eqs. (21) and (22)), the exact form of which is a matter of choice. A theory of pion photoproduction was constructed in the 1950s. Kroll and Ruderman [138] were the first to derive model-independent predictions in the threshold region, a so-called low energy theorem (LET), by applying gauge and Lorentz invariance to the reaction $\gamma N\to\pi N$. The general formalism for this process was developed by Chew, Goldberger, Low, and Nambu (Geoffrey Foucar Chew (1924-2018), Marvin Leonard “Murph” Goldberger (1922-2014), Francis Eugene Low (1921-2007), and Yoichiro Nambu (1921-2015) [^23] Nambu was awarded half of the Nobel Prize in Physics in 2008 for the discovery in 1960 of the mechanism of spontaneous broken symmetry in subatomic physics, related at first to the strong interaction’s chiral symmetry and later to the electroweak interaction and Higgs mechanism.). The results are known as CGLN amplitudes [139] and a few years later helicity amplitudes [140] were derived. About a decade later, Berends, Donnachie, and Weaver analyzed the existing data in terms of a multipole decomposition and extracted the various multipole amplitudes contributing in a region up to an excitation energy of $500~\mathrm{MeV}$ [141]. These amplitudes are vital inputs to low-energy descriptions of hadron physics based on chiral perturbation theory (ChPT) [142]. Within any of these bases, 16 possible bilinear combinations are referred to as the “observables.”
@@
Proton data alone do not allow for the separation of isoscalar and isovector components [143]. Accurate evaluation of EM couplings $N^{\ast}\to\gamma N$ and $\Delta\to\gamma N$ from meson photoproduction data remains a paramount task in hadron physics. Only with good data on both proton and neutron targets can one hope to disentangle isoscalar and isovector EM couplings of various $N^{\ast}$ and $\Delta^{\ast}$ resonances [144, 145], as well as the isospin properties of non-resonant background amplitudes.
@@
The lack of $\gamma n\to\pi^{-}p$ and $\gamma n\to\pi^{0}n$ data on free, isolated neutrons [^24] Apart from lower-energy ($700~\mathrm{MeV}$), there is data for the inverse $\pi^{-}$ photoproduction reaction, $\pi^{-}p\to\gamma n$ [112]. This process is free from complications associated with a deuteron target. However, there is a major disadvantage to using $\pi^{-}p\to\gamma n$: there is a large background from $\pi^{-}p\to\pi^{0}n\to\gamma\gamma n$, whose cross section is 5 to 500 times larger than $\pi^{-}p\to\gamma n$, which does not allow us to be as confident about the determination of neutron couplings relative to those of protons.
@@
Since the neutron targets do not exist, it remains necessary to use nuclear ones. In this case, when extracting information on the elementary reaction involving the bound neutron from nuclear data, one should take into account the nuclear-medium effects, i.e., the final-state interaction (FSI) and Fermi-motion effects [146, 147] [^25] It is impossible to measure FSI experimentally [148]. Obviously, in the case of polarized measurements, the FSI corrections are small or consistent with experimental uncertainties.
@@
Measurements of pion photoproduction on both proton and quasi-free neutron targets have a very long history, starting about 70 years ago with the discovery of the pion by the University of Bristol group [4]. Two years later, at the 1949 Spring Meeting of the US National Academy of Sciences, a preliminary account was given of some observations of mesons produced by the $335~\mathrm{MeV}$ photon beam from the Berkeley synchrotron [149] [^26] Edwin Mattison McMillan (1907-1991) shared the 1951 Nobel Prize in Chemistry with Glenn Theodore Seaborg (1912-1999), credited with being the first to produce a transuranium element, neptunium. In addition, McMillan co- invented the synchrotron with Vladimir Iosifovich Veksler (1907-1966). Starting with the use of bremsstrahlung facilities, pioneering results for $\gamma p\to\pi^{+}n$ (Fig. 14[149] and for $\gamma p\to\pi^{0}p$ [150] were obtained. One can possibly understand that the pion family is a triplet. Finding a neutral pion is much more difficult since it does not leave marks in photo-emulsion or Wilson chambers.
@@
In addition to the $\gamma p\to n\pi^{+}$ reaction on the free proton, measurements were performed at Berkeley on the loosely bound, quasi-free neutron in deuterium. Using a deuterium target, $\gamma n\to p\pi^{-}$ reactions were studied at the photon energy $318\pm 10~\mathrm{MeV}$ [151]. Hence, within three years, positive and negative pion photoproduction experiments were accomplished.
@@
Despite all the shortcomings of the first measurements (such as large normalization uncertainties, wide energy and angular binning, limited angular coverage, and so on), these data were crucial for the discovery of the first excited $\Delta$ and nucleon states using PWA for $\pi N$ elastic scattering data. The $\Delta(1232)3/2^{+}$ was determined by Enrico Fermi’s group [152]. While the second one $N(1440)1/2^{+}$, called Roper resonance, or colloquially simply “Roper”, in honor of L. David Roper (born 1935), who discovered this baryonic state, came several years after the $\Delta$-isobar [153] [^27] As an anecdote, it is worth mentioning that Lev Davidovich Landau, when becoming aware of Fermi’s discovery of the Delta resonance, did not believe in the existence of such a state, arguing A width of $120~\mathrm{MeV}$ - what is that? The pion will make a quarter of the circle around the nucleon, and that is supposed to be a pion-nucleon bound state? The concept of a meson-nucleon resonance was waiting to be established $\ldots$.
Figure 15: Database for $\gamma p\to\pi^{0}p$ (left) and $\gamma p\to\pi^{+}n$ (right). Experimental data from the SAID database [112] selected for 1996 through 2018. Right: Amount of data as a function of time. Full SAID database. The data is shown as a stacked histogram. Light shaded – cross sections, dark shaded – polarization data. The figure is adapted from Ref. [154].
Figure 16: Database for $\gamma n\to\pi^{-}p$ (left) and $\gamma n\to\pi^{0}n$ (right). Experimental data from the SAID database [112] selected for 1996 through 2018. Right: Amount of data as a function of time. Full SAID database. The data is shown as a stacked histogram. Light shaded – cross sections, dark shaded – polarization data. The figure is adapted from Ref. [154].
@@
The “neutron” database is significantly smaller than the “proton” one. The majority of single pseudoscalar meson photoproduction from the “neutron” target came from EM facilities at BNL and JLab, USA; MAMI, Germany; GRAAL, France (Fig. 16[154].
Studies of the $\gamma n\to\pi^{-}p$ and $\gamma n\to\pi^{0}n$ reactions can be carried out in quasi-free kinematics with deuteron targets. The reactions $\gamma d\to\pi^{-}p(p)$ and $\gamma d\to\pi^{0}n(p)$ in these kinematics involve a fast, knocked-out nucleon and a slow proton spectator; the slow proton is assumed not to be involved in the pion production process. In this quasi-free region, the reaction mechanism corresponds to the “dominant” impulse approximation (IA) diagram in Fig. 17(a) with the slow proton emerging from the deuteron vertex. Here, the differential cross section on the deuteron can be related to that on the neutron target in a well-understood way, see Ref. [155] and references therein. Fig. 17 illustrates this dominant IA diagram, as well as the leading terms of FSI corrections.
An energy- and angle-dependent FSI correction factor, R(E, $\theta$), can be defined as the ratio between the sum of three dominant diagrams in Fig. 17 and IA (the first of the diagrams). This can then be applied to the experimental $\gamma d$ data to obtain a two-body cross section for $\gamma n\to\pi^{-}p$ and $\gamma n\to\pi^{0}n$.
The GWU-ITEP FSI calculations (see Ref. [155] and references therein) are available over a broad energy range (threshold to $E=2.7~\mathrm{GeV}$), and for the full CM angular range ($\theta=0^{\circ}$ to $180^{\circ}$). Overall, the FSI correction factor $R<1.0$, while its value varies from 0.70 to 0.90 depending on the kinematics. The behavior of R is very smooth with respect to the pion production angle. There is a sizable FSI effect from the $S$-wave part of $pp$-FSI at small angles.
@@
R(E, $\theta$) is used as the FSI correction factor for the CLAS quasi-free $\gamma d\to\pi pN$ cross section averaged over the laboratory photon energy bin width. Note that the FSI correction grows rapidly in the forward direction ($\theta<30^{\circ}$). There are currently few measurements in this regime, so the uncertainty due to FSI for this reaction at forward angles does not cause much concern. The contribution of uncertainty in FSI calculations to the overall systematic normalization uncertainty is estimated to be about 2%–3% (the sensitivity to the deuteron wave function is 1%, and to the number of steps in the integration of the five-fold integrals is 2%. For the CLAS measurements, no sensitivity was found regarding the value of proton momentum used to determine whether or not it is a spectator.
@@
Figure 18: The elastic $\pi N$ partial wave scattering amplitudes in the isospin $I=3/2$ channel. The solid magenta lines and the green triangle points are the imaginary parts of the GiM and GWU/SAID amplitudes, respectively. The blue solid line and the red stars are the respective real parts of the GiM and GWU/SAID amplitudes.
Elastic Scattering of Pions on a Nucleon
Elastic scattering of hadrons is the process in which various kinds of interactions are reflected in a summarized manner. As emphasized in section 6, CC dynamics are much more involved than would be concluded from the elemental s-, t-, and u-channel meson-nucleon diagrams shown in Fig. 11. Induced interactions and the resulting polarization self-energies may be of paramount importance. In practice, however, their contributions will depend on the configuration space and, to some extent, on the general methodology of the approach. In order to identify and understand such dependencies, comparisons of model calculations are indispensable. In Fig. 18, the isospin $I=\frac{3}{2}$ partial wave scattering amplitudes obtained by the GiM approach (lines) and from the GWU/SAID project (symbols) are displayed for the entire spectrum of S-, P-, D-, and F-waves. Most prominently seen is the $P_{33}(1232)$ Delta resonance in the upper right panel, for which the two CC solutions indeed agree the best. Overall, the calculations are in fair agreement, albeit with detailed differences that are clearly visible, especially in the real parts of the amplitudes. They mainly reflect differences in the CC configuration space and different strategies for treating (or including at all) experimentally less confirmed resonances in the fitting procedures. The importance of such problems grows with energy and orbital angular momentum $L$. Overlapping and interfering structures complicate the clear identification of resonance structures.
3 etaeta and $eta^{\prime}$ Photoproduction on the Nucleon
@@
Overall, the experimental activities on the pseudoscalar-isoscalar $\eta(548)$ and $\eta^{\prime}(958)$ mesons have a short history, mainly because they are charge-neutral particles and, as such, are notoriously hard to detect. Compared to the pion case, the database for the isoscalar partner mesons is much smaller, as is evident from (Fig. 19).
In a pure quark-picture, one would assign $\eta\sim[q\bar{q}]$ as given by light $u,d$-quarks and $\eta^{\prime}\sim[s\bar{s}]$ as given by the heavier $s\bar{s}$-quarks. However, that assumption is in vain due to the remarkably large mass splitting of etaeta and $eta^{\prime}$ mesons by about $500~\mathrm{MeV}$. Moreover, the SU(3)-singlet $\eta^{\prime}$ meson is much heavier than the octet subset of the pseudoscalar nonet, which plays a central role in the spontaneous breaking of the $U(1)_{A}$ axial symmetry [156]. The mass splitting is caused by large mixing effects involving a bare gluonic flavor singlet $\eta_{0}$ state, as expressed by the famous Witten-Veneziano mass formula [157, 158]
$m^{2}_{\eta}+m^{2}_{\eta^{\prime}}=m^{2}_{K}+m^{2}_{\eta_{0}}$ (24)
which is satisfied for $m^{2}_{\eta_{0}}\simeq 0.71\mathrm{GeV}^{2}$. At Jefferson laboratory, the JEF (JLab Eta Factory) program aims to perform precision measurements of various $\eta$ and $\eta^{\prime}$ decay channels. The JEF collaboration expects especially clean data sets on the rare neutral decays of the two isoscalar mesons [159] .
@@
Figure 19: Database for $\gamma p\to\eta p$ (left) and $\gamma p\to\eta^{\prime}p$ (right). Experimental data from the SAID database [112] selected for 1996 through 2018. Right: Amount of data as a function of time. Full SAID database. The data is shown as a stacked histogram. Light shaded – cross sections, dark shaded – polarization data. The figure is adapted from Ref. [154].
Eta-production on the nucleon has a slightly different focus. Understanding the dynamics of $\eta$-meson production and, vice versa, the decay of nucleon resonances into the nucleon-eta exit channel is of ongoing interest in hadron spectroscopy. The $\eta$-meson photoproduction on the proton has been measured with high precision by the Crystal Ball Collaboration at MAMI [160]. These high-resolution data provide a new step forward in understanding the reaction dynamics and in the search for a signal from the “weak” resonance states. The main result reported in [160] is a very clean signal for a dip structure around $W=1.68~\mathrm{GeV}$, seemingly confirming older data [161, 162, 163, 164]. This raised the question of the origin of that structure, eventually indicating the appearance of a new, narrow, possibly exotic resonance state.
The study aimed to extend the previous coupled-channels analysis of the $\gamma p\to\eta p$ reaction by including the data from the new high-precision measurements [160]. The main question is whether the $\eta p$ reaction dynamics can be understood in terms of the established resonance states or whether a new state has to be introduced, thus confirming previous conjectures. A major issue for the analysis is unitarity and a consistent treatment of self-energy effects, as evidenced by the total decay width of resonances. Since the latter are driven by hadronic interactions, the analysis of photo-production data requires knowledge of the hadronic transition amplitudes as well. Hence, a coupled-channels description, such as the GiM, is an indispensable necessity.
As discussed in detail in Ref. [165], various relevant meson-baryon coupling constants were newly determined in the context of this work through large-scale coupled-channels calculations. This gave rise to improved constraints on the interaction parameters and the derived resonance parameters, i.e., masses and widths. Representative examples are the mass and width of the $D_{13}(1520)$ resonance, $M=1516\pm 10~\mathrm{MeV}$, and $\Gamma=106\pm 4~\mathrm{MeV}$, which agree with and confirm the values obtained earlier by Arndt et al. [166]. It is interesting to note that the mass of this resonance, deduced from pion photoproduction, tends to be $10~\mathrm{MeV}$ lower than the values derived from the pion-induced reactions [167]. The second $D_{13}(1900)$ state has a very large decay width. That state is likely to be related to the $D_{13}(2080)$ two-star resonance, proposed by the Particle Data Group in the 2024 issue of the Particle Physics Review [41].
The calculations demonstrate very satisfactory agreement with the experimental data in the entire kinematical region. The first peak is related to the $S_{11}(1535)$ resonance contribution. Similar to the $\pi^{-}p\to\eta n$ reaction, the $S_{11}(1650)$ and $S_{11}(1650)$ states interfere destructively, producing a dip around $W=1.68~\mathrm{GeV}$. The coherent sum of all partial waves leads to a more pronounced effect from the dip at forward angles.
In Fig. 21, results for the photon-beam asymmetry $\Sigma$ are compared to the GRAAL data. One can see that even close to the $\eta N$ threshold, where the calculations exhibit a dominant $S_{11}$ production mechanism, the beam asymmetry is non-vanishing for angles $\cos(\theta)\geq-0.2$. These results demonstrate that this observable is highly sensitive to very small contributions from higher partial waves. At $W=1.68~\mathrm{GeV}$ and forward angles, the GRAAL measurements show a rapid change in asymmetry behavior. In [168] this effect was explained by destructive interference between the $S_{11}(1535)$ and $S_{11}(1650)$ resonances, which induce the dip at $W\simeq 1.68~\mathrm{GeV}$ in the $S_{11}$ partial wave. Note that the interference between $S_{11}(1535)$ and $S_{11}(1650)$ and the interference between different partial waves are of a different nature. The overlapping of the $S_{11}(1535)$ and $S_{11}(1650)$ resonances do not simply mean a coherent sum of two independent contributions, but also include rescattering, i.e., coupled-channel effects. Such interplay can hardly be simulated by a simple sum of two Breit-Wigner distributions because this approach does not comply with unitarity.
4 Strangeness Production on the Nucleon by Kaon Photoproduction